OpenStudy (anonymous):

what is the expression in factored form x2-6x+8

OpenStudy (anonymous):

First list the factors of 8. They are: 1,2,4,8 Look at the factors of 8 and see what ones will give you the 6 that you need. 2 + 4 = 6 So we will use 2 7 4 to factor the equation. Notice that the six is negative while the eight is positive and that the six is larger than both of the factors that we want to use. This means that both factors are negative. so: x^2-6x+8=(x-2)(x-4) Check the answer (x-2)*(x-4) x^2-4x-2x+8----Combine like terms x^2-6x+8 The answer was right because we are back to where we started.